OpenStudy (anonymous):

write the ratio 5/9 to 7/9 as a fraction in lowest terms

OpenStudy (anonymous):

5/9 to 7/9 can be written as a fraction: \(\large \frac{\frac{5}{9}}{\frac{7}{9}}= \) ??? cancell the nines.... what u got then?

OpenStudy (anonymous):

5/7